# Discovery of Volatile Gas in the Giant Impact Disk around the 150-Myr old HD 23514

### v1

@@ Abstract
We report the discovery of CO2 gas emission around HD 23514, an F5V star in the 150 Myr-old Pleiades cluster, hosting one of the rare giant-impact disks with unique mineralogy dominated by silica dust. We show that the dust feature remains stable over several decades, and that the sub-µm grains, which give rise to the 9 µm feature, are co-spatial with the hot CO2 molecules within the sub-au vicinity of the star. Examining the Spitzer spectrum taken 15 years earlier, we show that the CO2 emission was also present at 4.3 significance. The existence of tiny silica grains and volatile gas requires special conditions to prevent the rapid loss caused by stellar radiation pressure and photodissociation. We explore several pathways explaining the observed properties and suggest that a past giant impact and/or stripping atmospheric event, involving large bodies with volatile content similar to the carbonaceous chondritic material, can simultaneously explain both the silica and volatile emission. Our discovery provides an important context for the amount of volatiles that a newly formed planet or the largest planetesimals could retain during the giant impact phase in the early solar system evolution.
@@ 1 Introduction
Young, extremely dusty disks, termed extreme debris disks (EDDs), give new insights into the collisional processes among planetary embryos during the era of terrestrial planet formation. These systems exhibit unique dust mineralogy dominated by thermodynamically altered minerals, likely produced by hypervelocity impacts, analogous to the Giant Impact Hypothesis for the formation of the Moon. The large amount of warm and highly processed dust definitively tie them to the rapid evolution of dust debris in their terrestrial zones, an environment traditionally thought to be dry and devoid of volatiles. Mid-infrared (mid-IR) monitoring shows that the majority of EDDs exhibit significant variability (Meng et al., 2014; Moór et al., 2021; Rieke et al., 2021; Su et al., 2023). In the best studied case of ID8 in NGC 2547, both yearly and monthly semi-periodic variability were observed, which was attributed to the collisional and dynamical evolution of a cloud of escaping boulders and vapor condensates formed in giant impact events (Meng et al., 2012; Su et al., 2019). Deep optical transits have also been found in some highly inclined systems where eclipsing dust clumps are found to be as large as the stars (de Wit et al., 2013; Gaidos et al., 2019; Melis et al., 2021; Su et al., 2022), further validating the scale of such collisions.
@@
HD 23514 (Cl* Melotte 22 HII 1132), an F5V star at a distance of 138.40.4 pc (Gaia Collaboration et al., 2021) in the Pleiades cluster, was first discovered to possess an exceptional amount of warm dust and unique dust mineralogy by Rhee et al. (2008). The age of the Pleiades, estimated to be 110–150 Myr (Lodieu et al., 2019), makes the system particularly interesting because it corresponds to the final building stage of terrestrial planet formation. After that stage, the rate of colossal Moon-forming events is expected to be significantly reduced (Quintana et al., 2016), consistent with that only 1% of the stars younger than a few hundred Myr show detectable 12 µm infrared excesses above 15% levels (Kennedy & Wyatt, 2013). To put HD 23514 in perspective, for 100 stars younger than the terrestrial planet formation stage, only one system is like HD 23514. Prior to the end of Spitzer cryogenic mission, Meng et al. (2012) showed that it is also one of the few systems exhibiting IR variability on timescales of a few years. The system shows no sign of cold (T40 K) dust and gas, suggesting the disk is compact, likely confined within 10 au (Vican et al., 2016; Sullivan et al., 2022). HD 23514 also has a wide M7 companion at a projected separation of 360 au and a position angle of 228° (Rodriguez et al., 2012), one of the characteristics that some EDDs share (Zuckerman, 2015; Moór et al., 2021, 2024).
@@
The HD 23514 system is one of the only three EDDs where the Spitzer/IRS spectra show prominent silica dust (Rhee et al., 2008; Lisse et al., 2009; Fujiwara et al., 2012), similar to those of the impact glasses and melted droplets commonly found in terrestrial impact craters (Morlok et al., 2016), therefore, they are also called giant impact disks. In the Giant Impact Hypothesis (Canup, 2004), the almost fully-grown Earth collided with a Mars-size planet called “Theia” at the end of its formation, and the material from the outer layers of both objects coalesced into the Moon. We can take this well-studied event as a prototype for such impacts. Because the Earth and Moon are almost chemically indistinguishable, Theia was thought to form at a similar distance from the Sun, i.e., dry material with significantly depleted volatile elements – those that vaporize easily upon heating. Some very young debris disks do contain gas, e.g., Pictoris (Roberge et al., 2000; Dent et al., 2014), but only typically at very young stages ( 50 Myr, Moór et al. 2017; Bonsor et al. 2023). Using JWST, we report the truly remarkable detection of hot molecular gas emission in HD 23514, particularly the CO2 lines have never been detected in debris disks before.
@@
Our observations and basic data reduction are described in Section 2. In Section 3, we describe the findings in terms of (1) lack of large-scale overall excess over four decades and the dust properties showing stable silica emission compared to the Spitzer data taken 15 years earlier, and (2) the properties of volatiles (robust detection of CO2 and CO, but tentative detections of H2O and NH3) and their likely existence in the Spitzer IRS data. In Section 4, we show that (1) the silica dust and gas volatiles are likely co-located within the sub-au region where their lifetimes are less than a year due to radiation pressure blowout and photodissociation, and (2) speculate on the origins of the stable silica dust and that of gas, in which both can be simultaneously explained by invoking a past giant impact event. The conclusion and future prospect are given in Section 5.
@@ 2 Observations and Data Reductions
HD 23514 was observed with MIRI/MRS on 2023 Sep 23, as part of the JWST GTO program PID1206 (PI: G. Rieke) to study the dust mineralogy and variability for a handful of EDDs discovered by Spitzer. All MRS channels and subbands were included, covering a wavelength range from 4.9 to 27.9 µm. Each subband was integrated with 12 groups in a single exposure using the FASTR1 mode with a total integration time of 132 s. A nominal observation setup was used, which includes the target acquisition with a neutral density filter, and a four-point dither pattern optimized for point sources.
We used the JWST Calibration Pipeline version 1.15 (Bushouse et al., 2024) and Calibration Reference Data System context pmap_1298 to reduce the data with default parameter setting and with additional steps for bad pixel self-calibration in the calwebb_spec2 and residual fringe correction in the calwebb_spec3 steps111see MIRI/MRS Pipeline Notebook https://github.com/spacetelescope/jwst-pipeline-notebooks/blob/main/notebooks/MIRI/MRS/JWPipeNB-MIRI-MRS.ipynb. The 12 spectral cubes were inspected visually to identify potential other sources within the IFU field of view. Only a very faint (by a few hundred times) object, 23 away from HD 23514 at a position angle of 148°, was found in the medium combined IFU images for the data shortward of 18 µm. This faint source is unlikely to be the late-M-type companion reported by Rodriguez et al. (2012) because of very different position angle and projected distance. Given its faintness, it has no impact in the extracted 1-D spectrum. We also examined the target spatial extension by comparing with a calibration star at all observed wavelengths, and concluded that the HD 23514 system is point-like and not resolved by JWST.
A 1D spectrum was extracted for each of the cubes using the default aperture and sky annulus (Law et al., 2025) that increased linearly with wavelength. Finally, to improve the continuity of the full spectrum, a small flux shift (a few %) was applied to individual spectra using the overlapping wavelength region and lining up with the shortest subband. Because the star contributes significant flux in the short wavelength part of the MIRI/MRS spectrum, we derive the disk emission by subtracting a model photosphere (details see Appendix A), yielding the spectrum shown in Figure 1 (left).
@@ 3.1 Lack of large-scale variation
As shown in Figure 1, the disk emission (in both flux level and solid-state feature) is remarkably similar to the one taken by Spitzer/IRS 15 years earlier. The prominent 9 µm silica feature agrees very well within 5% (1 uncertainty including the absolute flux calibration from both observatories), while in the broad 20 µm feature region the JWST data are consistently lower by 15% compared to the Spitzer one at 3 significance. This behavior suggests that there is no significant change in terms of dust composition, but the dust temperature is slightly lower in the JWST epoch (by 20 K, see Appendix A). The HD 23514 disk is known to display both weekly and yearly, stochastic IR variability, at 10–30% level in both the 3–5 µm and 24 µm photometry (Appendix B). Because there are no dust features in those bands, the variability is likely due to the change in the dust temperature rather than the composition. The overall level of the 10 µm feature has not changed appreciably over a 40-year baseline from the earliest IRAS 12 µm measurement taken in 1983, AKARI 9 µm taken in 2006, WISE 12 µm measurement taken in 2010 to the latest JWST data taken in 2023 (details see Figure 7). Lack of large-scale variation in the 10 µm feature over decades has also been reported for other EDDs like HD 113766 and HD 172555 (Su et al., 2020; Lisse et al., 2009; Samland et al., 2025).
@@ 3.2 Dust Properties
The mineral features are key in dust composition identifications due to stretching and bending motions in the silicon-oxygen anions. The narrow and strong resonance features from the thermally processed, crystalline silicates are easily distinguishable from the broad features of amorphous grains lacking ordered structure (Henning, 2010). The mass fraction of crystalline grains present in an infrared spectrum is then a good indication for the thermal processing in the disk environment (Sargent et al., 2006; Juhász et al., 2010). Watson et al. (2009) defined three crystalline mass fraction indices for the strongest features of pyroxenes (), olivines () and silica () in the 10 µm complex, centered at wavelengths 9.21, 11.08, and 12.46 µm, respectively, by referencing to a pristine (ISM-like, amorphous) silicate feature (for details, see Watson et al. 2009). An index value is roughly unity when the emission feature resembles the pristine profile, while an index value above unity reflects increasing prominence of a crystalline signature above the pristine one. The three indices are simple proxies for the amount of crystalline (thermally processed) grains without detailed dust decomposition modeling. Because the 9 µm silica peak can overlap with the 9.21 µm pyroxene peak under low spectral resolution, the silica index is defined by the narrow feature at 12.5 µm. In other words, the prominent 9 µm silica feature seen in HD 23514 is not entirely captured by the silica index (). Although these indices do not perfectly track the exact level of crystallinity on individual bases, they are good indicators when comparing properties among an ensemble of objects (Watson et al., 2009).
@@
Using the derived dust emissivity, the 10 µm emission complex in HD 23514 can be described as having a peak wavelength at 9.1 µm with a full-width-half-maximum (FWHM) of 2.04 µm, an equivalent width () of 4.70.2 µm, =2.54, =0.91, and =1.82. These crystalline indices suggest that the crystalline dust in HD 23514 is dominated by both pyroxene and silica with little of olivine type. The crystalline mass fraction is high (40% using the pyroxene index with the empirical trend derived by Watson et al. 2009), which is not common among PPDs. Silica dust has been found in a handful of PPDs, and one example (around a few Myr PPD, T51 (Sz 41)) is also shown in Figure 1 (right) for comparison. The 10 µm dust emissivity comparison between HD 23514 and the young star (Appendix A.2) suggests that the dust in HD 23514 has a higher crystallinity and the dominant dust size is smaller than that of the PPD (µm) as evidence by the feature’s FWHM. Furthermore, the (an indication for the amount of small grains) of 4.7 is also near the top among the 84 PPDs observed by Spitzer (only three systems have over 5 and all have near unity crystalline indices, i.e., low crystallinity fraction). The highly thermally processed dust mineralogy implies different formation processes. As detailed in Section 3.1 and Appendix A.2, past giant impact events are the most likely mechanism responsible for the large amount of highly processed crystalline silicates and silica grains in HD 23514.
@@ 3.3 Surprising Molecular Gas Emission Lines
In addition to the prominent silica feature, the high quality JWST spectrum also reveals surprising gas emission lines. Visual inspection indicates that the lines are dominated by CO2 emission centered at 14.98 µm (Q branch) accompanied by the P and R branches on the short- and long-wavelength sides. To properly characterize the molecular lines and discover other weaker ones, the dust continuum emission needs to be removed from the spectrum. The continuum was determined in an iterative way as detailed in Appendix C.1 and shown in the left panel of Figure 1. A subset of the continuum subtracted spectrum centered at 15 µm is shown in Figure 2 while the rest is shown in Figures 11 and 12.
@@
We used the local standard deviation per channel and subband as the typical root mean square (rms) of the continuum-subtracted MRS spectrum after masking out strong detected lines. Although the overall signal-to-noise ratio (S/N) of the data before continuum subtraction is 100 for 20 µm, unresolved emission lines with fluxes 5 mJy cannot be robustly detected in the current data. Despite the hints of H2O, and NH3 present in the spectrum, only CO2 Q-branch is detected at 10 along with the R and P branches at 3–5 , and CO at 2–4 . There is no sign of molecular/atomic hydrogen and hydrocarbon lines that are commonly found in PPDs, nor ionized/neutral atomic lines such as [Ne II], [Ar II], [Fe II], [Ni II], [Cl I], and [S I] present in HD 23514. Non-detection of molecular/atomic hydrogen and [Ne II] lines is the best indication that the gas in HD 23514 is not primordial but secondary (Pascucci et al., 2007), which is also consistent with the estimated upper limit for the CO-to-H2 mass ratio (Appendix C.2).
@@
To extract basic properties of the emitting gas, a plane-parallel slab model in local thermodynamic equilibrium (LTE) was adopted using three parameters: the gas emitting area (), column density (), and the gas temperature (). The emitting area is often expressed as an equivalent emitting radius () where . We focused on fitting the CO2 complex in the 13.5–16.3 µm region where the detected lines have the highest S/N. Assuming a turbulence velocity of zero, we find that a slab model with 0.0085 au2 (or 0.052 au), 1.38cm-2, and 891 K can reproduce the CO2 emission spectrum well. The associated uncertainties could be large (for details, see Appendix C.1), mostly due to the LTE assumption where the inferred temperature is higher, which could lead to an order magnitude lower column density than that in non-LTE (Bosman et al., 2017).
@@
Due to low S/N in most of the continuum-subtracted spectral region, we simply adopted similar parameters for H2O, NH3 and adjusted them to achieve reasonable fits visually. For CO, the same approach produces much narrower line widths in the limited wavelength range covered by MIRI/MRS, while an extra broadening along with a hotter gas temperature model can also reproduce the data (for details see Appendix C). Overall, the simple slab models can roughly match the peak flux of the CO lines in both cases, but the hotter one matches the line profiles better. We consider the detection of volatile CO2 and CO is robust, and the H2O and NH3 detections are tentative. Using the LTE slab model, the properties of CO2 are well constrained, but less so for CO and not at all for H2O and NH3. Non-LTE effects are particularly important for the low-density region that is very close to the star. Exploring non-LTE models is beyond the scope of the paper, which needs high S/N data (particularly for many of the weak water lines) and shorter wavelength coverage between 4.4 and 4.9 µm where the CO lines are brighter. Using LTE models we show that the JWST MIRI spectrum indubitably reveals the presence of hot (900 K) gas in the sub-au (0.03–0.05 au) region around HD 23514.

### v2

# Discovery of Volatile Gas in the Giant Impact Disk around the 150 Myr old HD 23514

@@ Abstract
We report the discovery of CO2 gas emission around HD 23514, an F5V star in the 150 Myr old Pleiades cluster, hosting one of the rare giant-impact disks with unique mineralogy dominated by silica dust. We show that the dust feature remains stable over several decades, and that the submicron grains, which give rise to the 9 µm feature, are cospatial with the hot CO2 molecules within the sub-au vicinity of the star. Examining the Spitzer spectrum taken 15 yr earlier, we show that the CO2 emission was also present at 4.3 significance. The existence of tiny silica grains and volatile gas requires special conditions to prevent the rapid loss caused by stellar radiation pressure and photodissociation. We explore several pathways explaining the observed properties and suggest that a past giant impact and/or stripping atmospheric event, involving large bodies with volatile content similar to the carbonaceous chondritic material, can simultaneously explain both the silica and volatile emission. Our discovery provides an important context for the amount of volatiles that a newly formed planet or the largest planetesimals could retain during the giant impact phase in the early solar system evolution.
@@ 1 Introduction
Young, extremely dusty disks, termed extreme debris disks (EDDs), give new insights into the collisional processes among planetary embryos during the era of terrestrial planet formation. These systems exhibit unique dust mineralogy dominated by thermodynamically altered minerals, likely produced by hypervelocity impacts, analogous to the giant-impact hypothesis for the formation of the Moon. The large amount of warm and highly processed dust definitively tie them to the rapid evolution of dust debris in their terrestrial zones, an environment traditionally thought to be dry and devoid of volatiles. Mid-infrared (mid-IR) monitoring shows that the majority of EDDs exhibit significant variability (Meng et al., 2014; Moór et al., 2021; Rieke et al., 2021; Su et al., 2023). In the best studied case of ID8 in NGC 2547, both yearly and monthly semiperiodic variability were observed, which were attributed to the collisional and dynamical evolution of a cloud of escaping boulders and vapor condensates formed in giant-impact events (Meng et al., 2012; Su et al., 2019). Deep optical transits have also been found in some highly inclined systems where eclipsing dust clumps are found to be as large as the stars (de Wit et al., 2013; Gaidos et al., 2019; Melis et al., 2021; Su et al., 2022), further validating the scale of such collisions.
@@
HD 23514 (Cl* Melotte 22 HII 1132), an F5V star at a distance of 138.40.4 pc (Gaia Collaboration et al., 2021) in the Pleiades cluster, was first discovered to possess an exceptional amount of warm dust and unique dust mineralogy by Rhee et al. (2008). The age of the Pleiades, estimated to be 110–150 Myr (Lodieu et al., 2019), makes the system particularly interesting because it corresponds to the final building stage of terrestrial planet formation. After that stage, the rate of colossal Moon-forming events is expected to be significantly reduced (Quintana et al., 2016), consistent with that only 1% of the stars younger than a few hundred million years show detectable 12 µm IR excesses above 15% levels (Kennedy & Wyatt, 2013). To put HD 23514 in perspective, for 100 stars younger than the terrestrial planet formation stage, only one system is like HD 23514. Prior to the end of Spitzer cryogenic mission, Meng et al. (2012) showed that it is also one of the few systems exhibiting IR variability on timescales of a few years. The system shows no sign of cold (T40 K) dust and gas, suggesting the disk is compact, likely confined within 10 au (Vican et al., 2016; Sullivan et al., 2022). HD 23514 also has a wide M7 companion at a projected separation of 360 au and a position angle of 228° (Rodriguez et al., 2012), one of the characteristics that some EDDs share (Zuckerman, 2015; Moór et al., 2021, 2024).
@@
The HD 23514 system is one of the only three EDDs where the Spitzer/IRS spectra show prominent silica dust (Rhee et al., 2008; Lisse et al., 2009; Fujiwara et al., 2012), similar to those of the impact glasses and melted droplets commonly found in terrestrial impact craters (Morlok et al., 2016), therefore, they are also called giant-impact disks. In the giant-impact hypothesis (Canup, 2004), the almost fully grown Earth collided with a Mars-size planet called Theia at the end of its formation, and the material from the outer layers of both objects coalesced into the Moon. We can take this well-studied event as a prototype for such impacts. Because the Earth and Moon are almost chemically indistinguishable, Theia was thought to form at a similar distance from the Sun, i.e., dry material with significantly depleted volatile elements – those that vaporize easily upon heating. Some very young debris disks do contain gas, e.g., Pic (Roberge et al., 2000; Dent et al., 2014), but only typically at very young stages ( 50 Myr, Moór et al. 2017; Bonsor et al. 2023). Using JWST, we report the truly remarkable detection of hot molecular gas emission in HD 23514, particularly the CO2 lines have never been detected in debris disks before.
@@
Our observations and basic data reduction are described in Section 2. In Section 3, we describe the findings in terms of (1) lack of large-scale overall excess over four decades and the dust properties showing stable silica emission compared to the Spitzer data taken 15 yr earlier, and (2) the properties of volatiles (robust detection of CO2 and CO, but tentative detections of H2O and NH3) and their likely existence in the Spitzer IRS data. In Section 4, we show that (1) the silica dust and gas volatiles are likely colocated within the sub-au region where their lifetimes are less than a year due to radiation-pressure blowout and photodissociation, and (2) speculate on the origins of the stable silica dust and that of gas, in which both can be simultaneously explained by invoking a past giant-impact event. The conclusion and future prospect are given in Section 5.
@@ 2 Observations and Data Reductions
HD 23514 was observed with MIRI/Medium-Resolution Spectrometer (MRS) on 2023 September 23, as part of the JWST GTO program PID1206 (PI: G. Rieke) to study the dust mineralogy and variability for a handful of EDDs discovered by Spitzer. All MRS channels and subbands were included, covering a wavelength range from 4.9 to 27.9 µm. Each subband was integrated with 12 groups in a single exposure using the FASTR1 mode with a total integration time of 132 s. A nominal observation setup was used, which includes the target acquisition with a neutral density filter, and a four-point dither pattern optimized for point sources.
We used the JWST Calibration Pipeline version 1.15 (Bushouse et al., 2024) and Calibration Reference Data System context pmap_1298 to reduce the data with default parameter settings and with additional steps for bad pixel self-calibration in the calwebb_spec2 and residual fringe correction in the calwebb_spec3 steps111see MIRI/MRS Pipeline Notebook at https://github.com/spacetelescope/jwst-pipeline-notebooks/blob/main/notebooks/MIRI/MRS/JWPipeNB-MIRI-MRS.ipynb. The 12 spectral cubes were inspected visually to identify potential other sources within the integral field unit (IFU) field of view. Only a very faint (by a few hundred times) object, 23 away from HD 23514 at a position angle of 148°, was found in the medium combined IFU images for the data shortward of 18 µm. This faint source is unlikely to be the late M-type companion reported by Rodriguez et al. (2012) because of very different position angle and projected distance. Given its faintness, it has no impact in the extracted 1D spectrum. We also examined the target spatial extension by comparing with a calibration star at all observed wavelengths, and concluded that the HD 23514 system is point-like and not resolved by JWST.
A 1D spectrum was extracted for each of the cubes using the default aperture and sky annulus (Law et al., 2025) that increased linearly with wavelength. Finally, to improve the continuity of the full spectrum, a small flux shift (a few percent) was applied to individual spectra using the overlapping wavelength region and lining up with the shortest subband. Because the star contributes significant flux in the short wavelength part of the MIRI/MRS spectrum, we derive the disk emission by subtracting a model photosphere (for details see Appendix A), yielding the spectrum shown in Figure 1 (left).
@@ 3.1 Lack of large-scale variation
As shown in Figure 1, the disk emission (in both flux level and solid-state feature) is remarkably similar to the one taken by Spitzer/IRS 15 yr earlier. The prominent 9 µm silica feature agrees very well within 5% (1 uncertainty including the absolute flux calibration from both observatories), while in the broad 20 µm feature region the JWST data are consistently lower by 15% compared to the Spitzer one at 3 significance. This behavior suggests that there is no significant change in terms of dust composition, but the dust temperature is slightly lower in the JWST epoch (by 20 K, see Appendix A). The HD 23514 disk is known to display both weekly and yearly stochastic IR variability, at 10–30% level in both the 3–5 µm and 24 µm photometry (Appendix B). Because there are no dust features in those bands, the variability is likely due to the change in the dust temperature rather than the composition. The overall level of the 10 µm feature has not changed appreciably over a 40 yr baseline from the earliest IRAS 12 µm measurement taken in 1983, AKARI 9 µm taken in 2006, WISE 12 µm measurement taken in 2010 to the latest JWST data taken in 2023 (details see Figure 7). Lack of large-scale variation in the 10 µm feature over decades has also been reported for other EDDs like HD 113766 and HD 172555 (Su et al., 2020; Lisse et al., 2009; Samland et al., 2025).
@@ 3.2 Dust Properties
Mineral features are key in dust composition identifications due to stretching and bending motions in the silicon-oxygen anions. The narrow and strong resonance features from the thermally processed, crystalline silicates are easily distinguishable from the broad features of amorphous grains lacking ordered structure (Henning, 2010). The mass fraction of crystalline grains present in an IR spectrum is then a good indication of thermal processing in the disk environment (Sargent et al., 2006; Juhász et al., 2010). Watson et al. (2009) defined three crystalline mass fraction indices for the strongest features of pyroxenes (), olivines () and silica () in the 10 µm complex, centered at wavelengths 9.21, 11.08, and 12.46 µm, respectively, by referencing to a pristine (ISM-like, amorphous) silicate feature (for details see Watson et al. 2009). An index value is roughly unity when the emission feature resembles the pristine profile, while an index value above unity reflects increasing prominence of a crystalline signature above the pristine one. The three indices are simple proxies for the amount of crystalline (thermally processed) grains without detailed dust decomposition modeling. Because the 9 µm silica peak can overlap with the 9.21 µm pyroxene peak under low spectral resolution, the silica index is defined by the narrow feature at 12.5 µm. In other words, the prominent 9 µm silica feature seen in HD 23514 is not entirely captured by the silica index (). Although these indices do not perfectly track the exact level of crystallinity on individual bases, they are good indicators when comparing properties among an ensemble of objects (Watson et al., 2009).
@@
Using the derived dust emissivity, the 10 µm emission complex in HD 23514 can be described as having a peak wavelength at 9.1 µm with a full-width-half-maximum (FWHM) of 2.04 µm, an equivalent width () of 4.70.2 µm, =2.54, =0.91, and =1.82. These crystalline indices suggest that the crystalline dust in HD 23514 is dominated by both pyroxene and silica with little of olivine type. The crystalline mass fraction is high (40% using the pyroxene index with the empirical trend derived by Watson et al. 2009), which is not common among PPDs. Silica dust has been found in a handful of PPDs, and one example (around a few megayear PPD, T51 (Sz 41)) is also shown in Figure 1 (right) for comparison. The 10 µm dust emissivity comparison between HD 23514 and the young star (Appendix A.2) suggests that the dust in HD 23514 has a higher crystallinity and the dominant dust size is smaller than that of the PPD (µm) as evidence by the feature’s FWHM. Furthermore, the (an indication of the amount of small grains) of 4.7 is also near the top among the 84 PPDs observed by Spitzer (only three systems have over 5 and all have near unity crystalline indices, i.e., low-crystallinity fraction). The highly thermally processed dust mineralogy implies different formation processes. As detailed in Section 3.1 and Appendix A.2, past giant-impact events are the most likely mechanism responsible for the large amount of highly processed crystalline silicates and silica grains in HD 23514.
@@ 3.3 Surprising Molecular Gas Emission Lines
In addition to the prominent silica feature, the high-quality JWST spectrum also reveals surprising gas emission lines. Visual inspection indicates that the lines are dominated by CO2 emission centered at 14.98 µm (Q branch) accompanied by the P and R branches on the short- and long-wavelength sides. To properly characterize the molecular lines and discover other weaker ones, the dust continuum emission needs to be removed from the spectrum. The continuum was determined in an iterative way as detailed in Appendix C.1 and shown in the left panel of Figure 1. A subset of the continuum-subtracted spectrum centered at 15 µm is shown in Figure 2 while the rest are shown in Figures 11 and 12.
@@
We used the local standard deviation per channel and subband as the typical rms of the continuum-subtracted MRS spectrum after masking out strong detected lines. Although the overall signal-to-noise ratio (S/N) of the data before continuum subtraction is 100 for 20 µm, unresolved emission lines with fluxes 5 mJy cannot be robustly detected in the current data. Despite the hints of H2O, and NH3 present in the spectrum, only CO2 Q-branch is detected at 10 along with the R and P branches at 3–5, and CO at 2–4. There is no sign of molecular/atomic hydrogen and hydrocarbon lines that are commonly found in PPDs, nor ionized/neutral atomic lines such as [Ne II], [Ar II], [Fe II], [Ni II], [Cl I], and [S I] present in HD 23514. Nondetection of molecular/atomic hydrogen and [Ne II] lines is the best indication that the gas in HD 23514 is not primordial but secondary (Pascucci et al., 2007), which is also consistent with the estimated upper limit for the CO-to-H2 mass ratio (Appendix C.2).
@@
To extract basic properties of the emitting gas, a plane-parallel slab model in local thermodynamic equilibrium (LTE) was adopted using three parameters: the gas emitting area (), column density (), and the gas temperature (). The emitting area is often expressed as an equivalent emitting radius () where . We focused on fitting the CO2 complex in the 13.5–16.3 µm region where the detected lines have the highest S/N. Assuming a turbulence velocity of zero, we find that a slab model with 0.0085 au2 (or 0.052 au), 1.38cm-2, and 891 K can reproduce the CO2 emission spectrum well. The associated uncertainties could be large (for details see Appendix C.1), mostly due to the LTE assumption where the inferred temperature is higher, which could lead to an order magnitude lower column density than that in non-LTE (Bosman et al., 2017).
@@
Due to low S/N in most of the continuum-subtracted spectral region, we simply adopted similar parameters for H2O, NH3 and adjusted them to achieve reasonable fits visually. For CO, the same approach produces much narrower line widths in the limited wavelength range covered by MIRI/MRS, while an extra broadening along with a hotter gas temperature model can also reproduce the data (for details see Appendix C). Overall, the simple slab models can roughly match the peak flux of the CO lines in both cases, but the hotter one matches the line profiles better. We consider the detections of volatile CO2 and CO is robust, and the H2O and NH3 detections are tentative. Using the LTE slab model, the properties of CO2 are well constrained, but less so for CO and not at all for H2O and NH3. Non-LTE effects are particularly important for the low-density region that is very close to the star. Exploring non-LTE models is beyond the scope of the paper, which needs high S/N data (particularly for many of the weak water lines) and shorter-wavelength coverage between 4.4 and 4.9 µm where the CO lines are brighter. Using LTE models we show that the JWST MIRI spectrum indubitably reveals the presence of hot (900 K) gas in the sub-au (0.03–0.05 au) region around HD 23514.